# Cataclysmic Variables in Triples: Formation Models and New Discoveries

### v1

@@ 1 Introduction
Cataclysmic Variables (CVs) are interacting binary stars in which a white dwarf accretes from a main-sequence (MS) companion star. Their orbital periods span from approximately min up to day, with a majority of systems found at periods of a few hours. (e.g., Patterson, 1984; Warner, 1995). To date, all evolutionary models of CVs have assumed that CVs form via common envelope (CE) evolution (e.g., Paczynski, 1976). In this framework, the primary star in the initially wide ( au) MS+MS progenitor binary evolves into a red giant (RG) and engulfs its companion. If the donor successfully ejects the envelope, the stars can emerge as a tight WD+MS binary. Subsequently, the MS star fills its Roche lobe, either due to angular momentum losses from magnetic braking and gravitational waves or as a result of its nuclear evolution, thereby forming the CV.
@@
The progenitors of most WDs in CVs have masses of a few M⊙ (e.g., Littlefair et al., 2008; Zorotovic & Schreiber, 2020). A notable fraction (; Evans et al., 2005; Raghavan et al., 2010; Tokovinin, 2014; Moe & Di Stefano, 2017; Offner et al., 2023) of such stars are actually triples, and this fraction can be even higher () considering that some of the wide third bodies have since become unbound (e.g., Shariat et al., 2023). Triple dynamics is known to significantly influence the evolution of other WD and MS binary populations (e.g., Naoz & Fabrycky, 2014; Toonen et al., 2016, 2017, 2020; Shariat et al., 2023, 2024b). It thus stands to reason that triple dynamics may play an important role in the formation of CVs. Yet, triples have been largely ignored in evolutionary models for CVs (with the notable exception of a triple model for the recurrent Nova T Pyx; Knigge et al., 2022), and there has never been a systematic observational study of CVs in triples.
@@
In hierarchical triples – where two stars are in a relatively close orbit (the inner binary) compared to the distant tertiary’s orbit around the inner binary – new dynamical phenomena are introduced, which alters the evolution of the inner binary. One major phenomenon in these systems is the eccentric Kozai Lidov (EKL) mechanism (Kozai, 1962; Lidov, 1962; Naoz, 2016), through which the tertiary can drive cyclic variations in the eccentricity and inclination of the inner binary in a secular fashion. For example, EKL-induced oscillations from a faraway tertiary can induce extreme eccentricities to the inner binary, leading to small ( R⊙) periastron separations. During these close passages, tides and mass transfer can shrink wide inner binaries and bring them into contact. This pathway can be a significant formation channel for many binary phenomena, such as black hole low-mass X-ray binaries (BH-LMXBs; Naoz et al., 2016; Shariat et al., 2024a), WD collisions (e.g., Thompson, 2011), nearly all close MS binaries (Tokovinin et al., 2006; Fabrycky & Tremaine, 2007), and several forms of WD or MS merger products (Lombardi et al., 2002; Naoz & Fabrycky, 2014; Cheng et al., 2019; Toonen et al., 2020; Heintz et al., 2022; Shariat et al., 2023, 2024b).
@@ 2.1 Constructing the Catalog
To identify a sample of cataclysmic variables (CVs) in triple systems, we cross-match three partially overlapping CV catalogs with a Gaia eDR3 wide binaries catalog. The first is the Ritter and Kolb (RK, 7th edition, last updated 2015) catalog, which includes CVs identified primarily through optical outbursts and X-ray associations (Ritter & Kolb, 2003). Unlike the following two catalogs, most of the CVs in RK have been followed up extensively and spectroscopically confirmed, leading to a high purity rate. The second source is the International Variable Star Index (VSX), from which we extract objects with CV classifications. Note that these classifications are often based on light curves and that some fractions are expected to be misclassified. The third catalog comprises CVs identified during outbursts via Zwicky Transient Facility (ZTF) transient alerts (Szkody et al., 2020, van Roestel et al. in prep). After consolidating these three catalogs and removing duplicates, we cross-match this initial sample with Gaia wide binaries. We search for wide stellar companions to the CVs by positionally cross-matching them to the wide binary catalog from El-Badry et al. (2021a). This catalog contains over a million spatially resolved binaries, where the two stars share a common proper motion and parallax in Gaia eDR3. The binary catalog is restricted to sources within kpc of the solar neighborhood, so we only consider CVs with 1 kpc.
El-Badry et al. (2021a) crudely classified the binaries in the catalog as ‘MSMS,’ ‘WDMS,’ and ‘WDWD’ based on the components’ position in the Gaia color-magnitude diagram, although there is expected contamination by unresolved inner binaries (e.g., El-Badry et al., 2021a; Nagarajan et al., 2024). Beyond MSMS binaries, which are the vast majority of the sample, the catalog also includes WDMS binaries, WDWD binaries, and binaries containing giants and subgiants. Most of the CVs in our sample fall between the MS and WD cooling track, with some classified as MS stars and others classified as WDs. The binary separations of the entire wide binary catalog range from a few au to around pc, with a peak near au that is mainly a consequence of Gaia’s arcsec angular resolution.
@@
=-0.5in Literature Name CV Catalog Spectroscopically Period Gaia DR3 ID [mas] [au] Confirmed? [hr] Gaia J051903.9+630339.6 285957277597658240 VSX, ZTF (1) – V3885 Sgr 6688624794231054976 RK, VSX (2) 4.97184 V379 Tel 6658737220627065984 RK, VSX, ZTF (1, 3) – V1108 Her 4538504384203931904 RK, VSX, ZTF (4) – MASTER OT J042609.34+354144.8 176429285061830144 RK, VSX, ZTF (5) – SDSS J101421.55+063857.7 3873721404734233344 ZTF (6) – LS IV -08 3 4339398736975240192 RK, VSX (7) 4.68695 ATO J145.8742+09.1629 3855044176807504768 VSX – 19.46712 YZ Cnc 683908812437792000 RK, VSX, ZTF (8) – USNO-A2.0 0825-10062737 4364235192818015744 VSX, ZTF – 1.272**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. KT Per 405873692214746368 RK, VSX (9) 3.90379 EF Peg 1759321791033449472 RK, VSX, ZTF (10) – UW Pic 4798833587650467200 RK, VSX (11) 2.22336 CSS 151208:020401+434133 346478795636689024 VSX, ZTF – 1.152**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. SDSS J120615.73+510047.0 1548430890981046272 RK (12) 3.792**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. ASASSN-19dn 5914778653652430464 VSX, ZTF – – RX J0131.4+3602 319016667370280064 ZTF – – ASASSN-21cw 4290033344968345088 VSX, ZTF – 8.544**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. RX J0154.0-5947 4714563374364671872 RK, VSX (13) 1.33440 MR UMa 772038105376131456 RK, VSX, ZTF (14) – IGR J19308+0530 4294249387962232576 VSX – 14.66208 Gaia DR3 2802565013608348928 2802565013608348928 ZTF – 7.776**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. DDE 202 4472788803304480896 VSX – 4.056**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. MASTER OT J072948.66+593824.4 1085880432371109888 RK, VSX, ZTF – 4.2**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. DDE 174 2203373473312011008 ZTF – – HW Boo 3741979368899159552 RK, VSX, ZTF (14) – DDE 169 377904796464796416 VSX, ZTF – – ASASSN-14hw 5549916268416336512 VSX, ZTF – – DDE 202 4580881692647497088 VSX, ZTF – 3.552**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. J2256+5954aaThis CV resides in a quadruple star system, where its wide companion is an unresolved binary. 2014349389931360768 RK (15) 5.48654 AH Men 5207384891323130368 RK, VSX (16) 3.05304 HS 0218+3229 325051822271077376 RK, VSX, ZTF (17) – ZTF18abmneqb 2861094289593259136 VSX, ZTF – 1.776**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. ASASSN-19pz 5391459325544320640 VSX, ZTF – – MASTER OT J084404.01+794408.7 1144277041110541440 VSX, ZTF – 2.52**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. BP CrA 6733050836430023552 VSX – – RZ Gru 6544371342567818496 RK, VSX (18) 8.64000 VZ Scl 2337436792938619392 RK, VSX (19) 3.47093 ASASSN-14eq 4918835764173746432 RK, VSX, ZTF (20) – NSV 7184 5985406470331309312 VSX, ZTF – – ASASSN-18ed 6139397368696634240 VSX, ZTF – – AY Psc 2565601982736199168 RK, VSX, ZTF (21) 5.232**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. SDSS J080033.86+192416.5 670132550216853632 RK, VSX, ZTF – – ASASSN-19wi 1777641510176281600 VSX, ZTF – – ASASSN-14jd 5269753451459457024 VSX, ZTF – – ASASSN-19eu 6099000383784197120 VSX, ZTF – – SWIFT J2124.6+0500 2699191408560964736 VSX (22) 19.992 TX Col 4804695423438691200 RK, VSX (23) 5.7192 References: (1) Pala et al. (2020); (2) Linnell et al. (2009); (3) Potter et al. (2005); (4) Price et al. (2004); (5) Han et al. (2020); (6) Inight et al. (2023); (7) Bruch & Diaz (2017); (8) Oke & Wade (1982); (9) Clarke & Bowyer (1984); (10) Howell & Liebert (1994); (11) Romero-Colmenero et al. (2003); (12) Schwope et al. (2009); (13) Beuermann et al. (2021); (14) Thorstensen & Skinner (2012); (15) Kjurkchieva et al. (2015); (16) Froning et al. (2012); (17) Rodríguez-Gil et al. (2009); (18) La Dous (1990); (19) Odonoghue et al. (1987); (20) Paterson et al. (2019); (21) Szkody & Howell (1993); (22) Halpern (2013); (23) Buckley & Tuohy (1987).
@@ 3.1 Numerical Setup
Our numerical simulations solve the three-body equations of motion up to the octupole level of approximation (full set of equations in Naoz, 2016). We also include stellar evolution for all three stars using the Single Stellar Evolution (SSE) code (Hurley et al., 2000). For both the inner and outer orbit, we include the effects of general relativistic precession to 1st post-Newtonian order (Naoz et al., 2013b), for both the inner and outer orbits. This reasonably models the dynamics of the mass ratios considered in this study (e.g., Naoz et al., 2013b; Lim & Rodriguez, 2020; Kuntz, 2022). For the inner binary, we also track tidal effects using an equilibrium tide model, allowing for the orbit to torque the stars and vice versa (e.g., Hut, 1980; Eggleton et al., 1998; Kiseleva et al., 1998; Naoz, 2016, and see latter appendix B from the full set of equations). Our tidal prescription changes based on the type of the star (Rose et al., 2019; Stephan et al., 2018, 2019, 2021). For main-sequence stars more massive than M⊙, we use a radiative tidal model. For low-mass main sequence stars ( M⊙) and red giants, we assume convective tides (Zahn, 1977). Additionally, the stars spin down due to magnetic braking, where the single-star spin-down rate is taken from (Zahn, 1977).
If the inner binary crosses its Roche limit (see below), we use COSMIC (Breivik et al., 2020) to model the evolution, which may include mass transfer. This effectively assumes that the inner binary is gravitationally decoupled from the tertiary during this time, which is reasonable given EKL is suppressed when tides dominate the apsidal precession of the inner binary (e.g., Antognini, 2015; Naoz, 2016). If the binary survives mass transfer and later becomes detached, we place it back into our triple code to continue its evolution. See Shariat et al. (2023) and Shariat et al. (2024b) for details on this triple evolution process. Simultaneously, we evolve the tertiary using single star tracks and follow the evolution of the outer orbit’s semi-major axis () and eccentricity () using adiabatic prescriptions (e.g., Shariat et al., 2023).
@@ 3.2 Initial Conditions of Three-Body Simulations
We draw initial orbital and stellar parameters from probability distributions similar to those used in Shariat et al. (2023). We draw from a Kroupa initial mass function (Kroupa et al., 1993) ranging from M⊙, where the range is chosen such that the primary will become a WD within the Gyr evolution time. We sample the inner binary mass ratio () from a uniform mass ratio distribution between to , and set (e.g., Raghavan et al., 2010; Moe & Di Stefano, 2017). We also sample the mass ratio of the outer binary () from the same uniform distribution and set . Previous triple population synthesis studies show that population-level results are relatively insensitive to the choice of mass ratio distribution (Shariat et al., 2023). The initial mutual inclinations are chosen from an isotropic distribution (uniform in ), and the initial and are chosen from a uniform distribution, as are the initial spin-orbit angles for the stars in the inner binary (). The periods of the inner and outer orbit ( and ) are drawn from Raghavan et al. (2010) distribution, which is a Gaussian in with and .
@@
A deviation from hierarchy does not necessarily mean that a triple becomes unstable immediately (Grishin et al., 2017; Mushkin & Katz, 2020; Bhaskar et al., 2021; Toonen et al., 2022; Zhang et al., 2023). However, given that all of the CV triples in our observational sample are in the very hierarchical regime, we require hierarchy in our simulations based on Equation (3).
Figure 5 depicts the initial condition distributions in this study. We show the ditribution of initial masses (, , ), semi-major axes (, ), eccentricities (, ), and mutual inclination (). On the top of Figure 5, we show a schematic representation of a CV in a hierarchical triple (not to scale) along with some of the key parameters labeled (masses, eccentricities, and semi-major axes).
We run each triple simulation for an upper limit of Gyr, but systems are stopped earlier if the inner binary crosses the Roche limit or merges entirely.

### v2

@@ 1 Introduction
Cataclysmic Variables (CVs) are interacting binary systems in which a white dwarf accretes from a main-sequence or partially evolved companion star. Their orbital periods span from approximately min up to day, with a majority of systems found at periods of a few hours. (e.g., Patterson, 1984; Warner, 1995). To date, all evolutionary models of CVs (outside of dense clusters) have assumed that CVs form via common envelope (CE) evolution (e.g., Paczynski, 1976). In this framework, the primary star in the initially wide ( au) MS+MS progenitor binary evolves into a red giant (RG) and engulfs its companion. If the donor successfully ejects the envelope, the stars can emerge as a tight WD+MS binary. Subsequently, the MS star fills its Roche lobe, either due to angular momentum losses from magnetic braking and gravitational waves or as a result of its nuclear evolution, thereby forming the CV.
@@
The progenitors of most WDs in CVs have masses of a few M⊙ (e.g., Littlefair et al., 2008; Zorotovic & Schreiber, 2020). A notable fraction (; Evans et al., 2005; Raghavan et al., 2010; Tokovinin, 2014; Moe & Di Stefano, 2017; Offner et al., 2023) of such stars are actually triples, and this fraction can be even higher () considering that some of the wide third bodies have since become unbound (e.g., Shariat et al., 2023). Triple dynamics is known to significantly influence the evolution of other WD and MS binary populations (e.g., Naoz & Fabrycky, 2014; Toonen et al., 2016, 2017, 2020; Shariat et al., 2023, 2024). It thus stands to reason that triple dynamics may play an important role in the formation of CVs. Yet, triples have been largely ignored in evolutionary models for CVs (with the notable exception of a triple model for the recurrent Nova T Pyx; Knigge et al., 2022), and there has never been a systematic observational study of CVs in triples.
@@
In hierarchical triples – where two stars are in a relatively close orbit (the inner binary) compared to the distant tertiary’s orbit around the inner binary – new dynamical phenomena are introduced, which alters the evolution of the inner binary. One major phenomenon in these systems is the eccentric Kozai Lidov (EKL) mechanism (Kozai, 1962; Lidov, 1962; Naoz, 2016), through which the tertiary can drive cyclic variations in the eccentricity and inclination of the inner binary in a secular fashion. For example, EKL-induced oscillations from a faraway tertiary can induce extreme eccentricities to the inner binary, leading to small ( R⊙) periastron separations. During these close passages, tides and mass transfer can shrink wide inner binaries and bring them into contact. This pathway can be a significant formation channel for many binary phenomena, such as black hole low-mass X-ray binaries (BH-LMXBs; Naoz et al., 2016; Shariat et al., 2025), WD collisions (e.g., Thompson, 2011), nearly all close MS binaries (Tokovinin et al., 2006; Fabrycky & Tremaine, 2007), and several forms of WD or MS merger products (Lombardi et al., 2002; Naoz & Fabrycky, 2014; Cheng et al., 2019; Toonen et al., 2020; Heintz et al., 2022; Shariat et al., 2023, 2024).
@@ 2.1 Constructing the Catalog
To identify a sample of cataclysmic variables (CVs) in triple systems, we cross-match three partially overlapping CV catalogs with a Gaia eDR3 wide binaries catalog. The first is the Ritter and Kolb (RK, 7th edition, last updated 2015) catalog, which includes CVs identified primarily through optical outbursts and X-ray associations (Ritter & Kolb, 2003). Unlike the following two catalogs, most of the CVs in RK have been followed up extensively and spectroscopically confirmed, leading to a high purity rate. The second source is the International Variable Star Index (VSX), from which we extract objects with CV classifications. Note that these classifications are often based on light curves and that some fractions are expected to be misclassified. The third catalog comprises CVs identified during outbursts via Zwicky Transient Facility (ZTF) transient alerts (Szkody et al., 2020, van Roestel et al. in prep). After consolidating these three catalogs and removing duplicates, we cross-match this initial sample with Gaia wide binaries. We search for wide stellar companions to the CVs by positionally cross-matching them to the wide binary catalog from El-Badry et al. (2021a). This catalog contains over a million spatially resolved binaries, where the two stars share a common proper motion and parallax in Gaia eDR3. The binary catalog is restricted to sources within kpc of the Sun, so we only consider CVs with 1 kpc. Previous works (Michaely & Perets, 2014; Igoshev et al., 2020) proposed and carried out searches for common proper motion companions around post-CE binaries to constrain common-envelope evolution. Here, we adopt a similar methodology to study formation pathways of CVs, using the more precise Gaia DR3 astrometry and incorporating explicit chance alignment probabilities to identify wide companions with high confidence.
El-Badry et al. (2021a) crudely classified the binaries in the catalog as ‘MSMS,’ ‘WDMS,’ and ‘WDWD’ based on the components’ position in the Gaia color-magnitude diagram, although there is expected contamination by unresolved inner binaries (e.g., El-Badry et al., 2021a; Nagarajan & El-Badry, 2024). Beyond MSMS binaries, which are the vast majority of the sample, the catalog also includes WDMS binaries, WDWD binaries, and binaries containing giants and subgiants. Most of the CVs in our sample fall between the MS and WD cooling track, with some classified as MS stars and others classified as WDs. The binary separations of the entire wide binary catalog range from a few au to around pc, with a peak near au that is mainly a consequence of Gaia’s arcsec angular resolution.
@@
=-0.5in Literature Name CV Catalog Spectroscopically Period Gaia DR3 ID [mas] [au] Confirmed? [hr] Gaia J051903.9+630339.6 285957277597658240 VSX, ZTF (1) – V3885 Sgr 6688624794231054976 RK, VSX (2) 4.97184 V379 Tel 6658737220627065984 RK, VSX, ZTF (1, 3) – V1108 Her 4538504384210935424 RK, VSX, ZTF (4) – MASTER OT J042609.34+354144.8 176429285061830144 RK, VSX, ZTF (5) – SDSS J101421.55+063857.7 3873721404734233344 ZTF (6) – LS IV -08 3 4339398736975240192 RK, VSX (7) 4.68695 ATO J145.8742+09.1629 3855044176807504768 VSX – 19.46712 YZ Cnc 683908812437792000 RK, VSX, ZTF (8) – USNO-A2.0 0825-10062737 4364235192818015744 VSX, ZTF – 1.272**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. KT Per 405873692214746368 RK, VSX (9) 3.90379 EF Peg 1759321791033449472 RK, VSX, ZTF (10) – UW Pic 4798833587650467200 RK, VSX (11) 2.22336 CSS 151208:020401+434133 346478795636689024 VSX, ZTF – 1.152**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. SDSS J120615.73+510047.0 1548430890981046272 RK (12) 3.792**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. ASASSN-19dn 5914778653652430464 VSX, ZTF – – RX J0131.4+3602 319016667370280064 ZTF – – ASASSN-21cw 4290033344968345088 VSX, ZTF – 8.544**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. RX J0154.0-5947 4714563374364671872 RK, VSX (13) 1.33440 MR UMa 772038105376131456 RK, VSX, ZTF (14) – IGR J19308+0530 4294249387962232576 VSX – 14.66208 Gaia DR3 2802565013608348928 2802565013608348928 ZTF – 7.776**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. DDE 202 4472788803304480896 VSX – 4.056**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. MASTER OT J072948.66+593824.4 1085880432371109888 RK, VSX, ZTF – 4.2**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. DDE 174 2203373473312011008 ZTF – – HW Boo 3741979368899159552 RK, VSX, ZTF (14) – DDE 169 377904796464796416 VSX, ZTF – – ASASSN-14hw 5549916268416336512 VSX, ZTF – – DDE 202 4580881692647497088 VSX, ZTF – 3.552**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. J2256+5954aaThis CV resides in a quadruple star system, where its wide companion is an unresolved binary. 2014349389931360768 RK (15) 5.48654 AH Men 5207384891323130368 RK, VSX (16) 3.05304 HS 0218+3229 325051822271077376 RK, VSX, ZTF (17) – ZTF18abmneqb 2861094289593259136 VSX, ZTF – 1.776**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. ASASSN-19pz 5391459325544320640 VSX, ZTF – – MASTER OT J084404.01+794408.7 1144277041110541440 VSX, ZTF – 2.52**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. BP CrA 6733050836430023552 VSX – – RZ Gru 6544371342567818496 RK, VSX (18) 8.64000 VZ Scl 2337436792938619392 RK, VSX (19) 3.47093 ASASSN-14eq 4918835764173746432 RK, VSX, ZTF (20) – NSV 7184 5985406470331309312 VSX, ZTF – – ASASSN-18ed 6139397368696634240 VSX, ZTF – – AY Psc 2565601982736199168 RK, VSX, ZTF (21) 5.232**Denotes the best-fit period from the ZTF optical light curve for sources without reported VSX periods. SDSS J080033.86+192416.5 670132550216853632 RK, VSX, ZTF – – ASASSN-19wi 1777641510176281600 VSX, ZTF – – ASASSN-14jd 5269753451459457024 VSX, ZTF – – ASASSN-19eu 6099000383784197120 VSX, ZTF – – SWIFT J2124.6+0500 2699191408560964736 VSX (22) 19.992 TX Col 4804695423438691200 RK, VSX (23) 5.7192 References: (1) Pala et al. (2020); (2) Linnell et al. (2009); (3) Potter et al. (2005); (4) Price et al. (2004); (5) Han et al. (2020); (6) Inight et al. (2023); (7) Bruch & Diaz (2017); (8) Oke & Wade (1982); (9) Clarke & Bowyer (1984); (10) Howell & Liebert (1994); (11) Romero-Colmenero et al. (2003); (12) Schwope et al. (2009); (13) Beuermann et al. (2021); (14) Thorstensen & Skinner (2012); (15) Kjurkchieva et al. (2015); (16) Froning et al. (2012); (17) Rodríguez-Gil et al. (2009); (18) La Dous (1990); (19) Odonoghue et al. (1987); (20) Paterson et al. (2019); (21) Szkody & Howell (1993); (22) Halpern (2013); (23) Buckley & Tuohy (1987).
@@ 3.1 Numerical Setup
Our numerical simulations solve the three-body equations of motion up to the octupole level of approximation (full set of equations in Naoz, 2016). We also include stellar evolution for all three stars using the Single Stellar Evolution (SSE) code (Hurley et al., 2000). For both the inner and outer orbit, we include the effects of general relativistic precession to 1st post-Newtonian order (Naoz et al., 2013b). This reasonably models the dynamics of the mass ratios considered in this study (e.g., Naoz et al., 2013b; Lim & Rodriguez, 2020; Kuntz, 2022). For the inner binary, we also track tidal effects using an equilibrium tide model, allowing for the orbit to torque the stars and vice versa (e.g., Hut, 1980; Eggleton et al., 1998; Kiseleva et al., 1998; Naoz, 2016, and see latter appendix B from the full set of equations). Our tidal prescription changes based on the type of the star (Rose et al., 2019; Stephan et al., 2018, 2019, 2021). For main-sequence stars more massive than M⊙, we use a radiative tidal model. For low-mass main sequence stars ( M⊙) and red giants, we assume convective tides (Zahn, 1977). Additionally, the stars spin down due to magnetic braking, where the single-star spin-down rate is taken from (Zahn, 1977).
If the inner binary crosses its Roche limit (see below), we use COSMIC (Breivik et al., 2020) to model the evolution, which may include mass transfer. This effectively assumes that the inner binary is gravitationally decoupled from the tertiary during this time, which is reasonable given EKL is suppressed when tides dominate the apsidal precession of the inner binary (e.g., Antognini, 2015; Naoz, 2016). If the binary survives mass transfer and later becomes detached, we place it back into our triple code to continue its evolution. See Shariat et al. (2023) and Shariat et al. (2024) for details on this triple evolution process. Simultaneously, we evolve the tertiary using single star tracks and follow the evolution of the outer orbit’s semi-major axis () and eccentricity () using adiabatic prescriptions (similar to Shariat et al., 2023).
@@ 3.2 Initial Conditions of Three-Body Simulations
We draw initial orbital and stellar parameters from probability distributions similar to those used in Shariat et al. (2023). We draw from a Kroupa initial mass function (Kroupa et al., 1993) ranging from M⊙, where the range is chosen such that the primary will become a WD within the Gyr evolution time. We sample the inner binary mass ratio () from a uniform mass ratio distribution between to , and set (e.g., Raghavan et al., 2010; Moe & Di Stefano, 2017). We also sample the mass ratio of the outer binary () from the same uniform distribution and set . Previous triple population synthesis studies show that population-level results are relatively insensitive to the choice of mass ratio distribution (Shariat et al., 2023). The initial mutual inclinations are chosen from an isotropic distribution (uniform in ). Observed triples with similar separations and component masses to our simulated systems show a median mutual inclination of (Tokovinin, 2022), consistent with both isotropic and uniform distributions444 In contrast, compact hierarchical triples ( days) exhibit a coplanar excess (e.g., Bashi & Tokovinin, 2024), though these are not the focus of this study.. The initial and are chosen from a uniform distribution, as are the initial spin-orbit angles for the stars in the inner binary (). The periods of the inner and outer orbit ( and ) are drawn from Raghavan et al. (2010) distribution, which is a Gaussian in with and .
@@
A deviation from hierarchy does not necessarily mean that a triple becomes unstable immediately (Grishin et al., 2017; Mushkin & Katz, 2020; Bhaskar et al., 2021; Toonen et al., 2022; Zhang et al., 2023). However, given that all of the CV triples in our observational sample are in the very hierarchical regime, we require hierarchy in our simulations based on Equation (3)555 We also evaluate the stability of our systems using the more recent stability criterion from Vynatheya et al. (2022), finding find that most () are stable under their algebraic framework. In mildly hierarchical triples, quasi-secular dynamics can influence orbital evolution by breaking down the assumptions of standard, double-orbit-averaged secular theory (e.g., Antonini & Perets, 2012; Katz & Dong, 2012; Antonini et al., 2014; Luo et al., 2016; Grishin et al., 2018; Tremaine, 2023; Klein & Katz, 2024). However, in our simulated sample, the vast majority of systems () lie in the extremely hierarchical regime, with EKL timescales much longer than the outer orbital period (median = 0.005). This suggests that quasi-secular effects are unlikely to play a significant role in the evolution of most systems in our study. We run each triple simulation for an upper limit of Gyr, but systems are stopped earlier if the inner binary crosses the Roche limit or merges entirely.
Figure 5 depicts the initial condition distributions in this study. We show the distribution of initial masses (, , ), semi-major axes (, ), eccentricities (, ), and mutual inclination (). On the top of Figure 5, we show a schematic representation of a CV in a hierarchical triple (not to scale) along with some of the key parameters labeled (masses, eccentricities, and semi-major axes).